If a firm in perfect competition is producing at the quantity where MR = MC why will it not produce any more?
Answer #1

We know that profit of a firm is givedn by \pi = Total revenue(TR)- Total cost(C). Now to maximise profit, the profit maximising condition will be: d\pi/dQ=0

Now we know that total revenue= PQ where P= price level and Q= output level. So,

d\pi/dQ= d(TR)/dQ - dC/dQ = 0

Now, d(TR)/dQ= marginal revenue= MR and dC/dQ= marginal cost. Hence d\pi/dQ=0 implies that MR-MC=0

Or, MR=MC. Hence in case of perfect competition profit is maximised when marginal revenue becomes equal to marginal cost and beyond this level, MC becomes greater than MR or the increment in cost by producing an extra unit becomes greater than the increment in revenue by producing an extra unit. Hence it will not be profitable for a perfectly competitive firm to produce beyond the MR=MC level.